jchristiansen:

Hannah’s favorite toy character is sold wearing one of many outfits. An outfit consists of one shirt, one pair of pants, and one cap. The shirt can be striped, plain, or polka-dotted. The pants may be denim or plaid. The cap may be a cowboy hat or a baseball cap. If Hannah picks a toy at random from the shelf and there are an equal number of toys wearing each outfit, what is the probability that the toy is dressed in a polka-dotted shirt and a cowboy hat?

justjm:

This question revolves around the idea of combinatorics and multiplication rule. P(Poka-dotted) = 1/3 P(Cowboy hat) = 1/2 P(Poka-dotted AND Cowboy hat) = P(Poka-dotted)P(Cowboy Hat)